OpenStudy (anonymous):

how to write 2000 as a product of primes

OpenStudy (kc_kennylau):

\[\large2000=2\times1000=2\times10^3=2\times(2\times5)^3=2^4\times5^3\]

OpenStudy (kc_kennylau):

(That was what goes on your mind, here's what you write) \[\large2000=2\times1000=2^2\times500=2^3\times250=2^4\times125\\\large\hspace{26pt}=2^4\times5\times25=2^4\times5^2\times5=2^4\times5^3\]